OpenStudy (lynda.boyce):

I am having issues graphing a problem. T '(r)= -ln1.50/12(1/(10000/833 +r)(ln(1+0.0833r))^2) can anyone help?

OpenStudy (anonymous):

You would like to graph this?

OpenStudy (lynda.boyce):

yes, please! I have been trying to go to different sites to graph it but I just can't figure out how to do it

OpenStudy (anonymous):

can you post the original problem, how you derived this expression

OpenStudy (anonymous):

It looks like a derivative?

OpenStudy (lynda.boyce):

My k value I chose was 1.50.

OpenStudy (anonymous):

and you want to graph the derivative of this

OpenStudy (lynda.boyce):

I am to graph the T(r) and the T'(r). I have already graphed the T(r). I am having issues with the T'(r)

OpenStudy (anonymous):

Here is the graph of the derivative https://www.desmos.com/calculator/ntdrf9yodq

OpenStudy (lynda.boyce):

Thank you so much!

OpenStudy (anonymous):

to get the derivative f(r) = ln(1.5)/(12*ln(1+0.0833*r)) first rewrite the expression f(r) = ln(1.5)/12 * [ ln(1+0.0833*r) ]^-1 f ' (r) = ln(1.5)/12 * (-1)[ ln(1+0.0833*r) ]^-2 * (0.0833) \[f'(r) = \frac{ (-\ln(1.5)*0.0833/12)}{ (\ln(1+0.0833*r)) ^2} \]